# Constraining $\boldsymbol{|V_{cs}|}$ and physics beyond the Standard Model from exclusive (semi)leptonic charm decays

### v1

@@ 3 Methods and Results
nominal SM $60.3$ $51$ $17.5$ $239.1\pm 0.3$
CKM $51.9$ $50$ $40.0$ $251.4\pm 0.3$
WET $47.7$ $42$ $25.1$ $251.1\pm 0.3$
scale factor SM $67.4$ $51$ $6.1$ $231.5\pm 0.3$
CKM $47.3$ $50$ $58.3$ $248.8\pm 0.3$
WET $56.5$ $42$ $6.7$ $252.6\pm 0.3$
@@
Figure 2: Comparison of the predictions in our “nominal” scenario for the various branching ratios for $D_{s}^{+}\to\ell^{+}\nu$, $D^{+(0)}\to K^{0}_{S}(K^{-})\ell^{+}\nu$, and $\Lambda_{c}\to\Lambda\ell^{+}\nu$. The decay $D_{s}^{*+}\to\ell^{+}\nu$ is omitted, since the large theory uncertainties make a meaningful visual comparison off the predictions with the measurement impossible. Measurements are shown as black diamonds. Prior predictions labelled as “theory only” are shown as purple coloured circles. Posterior predictions for the two fit models are shown as magenta triangles (SM) and cyan squares (CKM).
@@ 3.1 SM Prior Predictions and Fit
To test a possible issue with the normalisation of the lattice QCD results, we float all hadronic nuisance parameters within the SM fit model (“nominal” scenario) and fit to the full experimental likelihood. We find that we can reduce the experimental $\chi^{2}$ to $60.3$ with a $p$ value of $17.5\%$. This happens at the expense of moving away from the a-priori parameter values by $4.9\,\sigma$. We produce posterior-predictive distributions for the relevant absolute branching fractions, which are shown in fig. 2 as magenta triangles. We can still observe that the posterior predictions systematically overshoot the measurements, however, the tensions with respect to the experimental measurements are visibly reduced. At the level of the variety, number, and accuracy of the hadronic matrix elements used in this analysis, we find compatibility between the hadronic matrix elements and the dispersive bounds.
@@ 3.2 CKM Fit
which agrees with the weighted average of the individual fit results. The central value for the “scale factor” result is higher than the above but compatible at the $1\,\sigma$ level, as shown on the right-hand side of fig. 3. As illustrated in fig. 2, floating the CKM parameter $|V_{cs}|$ results in overall better agreement with the data. Contrary to the SM prior predictions and the SM fit posterior predictions, no systematic shift to larger or smaller values for the branching ratios are visible. This is also reflected in table 1. For the “nominal” scenario, we show that our CKM fit yields a reduction of the $\chi^{2}$ in the best-fit point by $9$ at the expense of one degree of freedom. Using Wilks’ theorem, we therefore obtain a preference of the CKM fit over the SM fit at the $2.7\,\sigma$ level. A Bayesian model comparison yields a Bayes factor of $\simeq 2.4\cdot 10^{4}$, decisively in favour of the CKM fit model. For the “scale factor” scenario we find a similar preference for the CKM fit model.
@@ 3.3 CKM unitarity
The violation of CKM unitarity of around $3\,\sigma$ by the first-row elements of the CKM matrix has received quite some attention Seng:2018yzq ; Belfatto:2019swo ; Grossman:2019bzp ; Crivellin:2020lzu ; Kirk:2020wdk ; Crivellin:2021njn ; Seng:2021nar ; Cirigliano:2022yyo ; Crivellin:2022rhw . Here we perform an alternative test of CKM unitarity by probing the normalisation of both the second row and column;
@@
Our results are given in table 3. Under the assumption of $100\%$ positively correlated uncertainties, we find a deficit toward the expectation of CKM second row and second-column unitarity at the $4.3\,\sigma$ level and $5.2\,\sigma$ level using the nominal results.
@@
Table 3: Results of the CKM unitarity tests. The last two lines contain the squared sum of CKM element of second row or column, as well as the pull to unity assuming Gaussian uncertainties. The uncertainty on the CKM entries are assumed 100% positively correlated, which corresponds to the most conservative scenario (proper estimations would give larger pulls).
This large tension with unitarity again strengthens the case to investigate the normalisation issue that we already point out in section 3.1.
@@ 3.4 BSM Interpretation
Our findings so far motivate to further investigate the allowed parameter space for BSM contribution to $c\to s\ell^{+}\nu$ processes. Lifting the assumption of SM dynamics, we fit the 9 parameters discussed in section 2.4. We find the resulting posterior PDF to be multi modal and each mode to be distinctly non-Gaussian. The individual modes are related through symmetries of our likelihood and therefore feature the same maximum a-posteriori. Here and in our supplementary material, we provide information on one chosen mode of the posterior, which is defined by $\arg\mathcal{C}^{\ell}_{V,L}=0$ and $\mathcal{C}^{\ell}_{V,L}\simeq 1$. In our nominal scenario, we obtain the following 68% probability intervals
$\operatorname{Re}\mathcal{C}^{\ell}_{V,L}$ $=[\phantom{-}0.941,\phantom{-}0.998]\,,$ (27)
$\operatorname{Re}\mathcal{C}^{\ell}_{V,R}$ $=[-0.023,-0.009]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{V,R}$ $=[-0.277,0.277]\,,$
$\operatorname{Re}\mathcal{C}^{\ell}_{S,L}$ $=[-0.018,\phantom{-}0.015]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{S,L}$ $=[-0.028,0.028]\,,$
$\operatorname{Re}\mathcal{C}^{\ell}_{S,R}$ $=[-0.024,\phantom{-}0.009]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{S,R}$ $=[-0.029,0.029]\,,$
$\operatorname{Re}\mathcal{C}^{\ell}_{T}$ $=[-0.023,\phantom{-}0.045]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{T}$ $=[-0.065,0.065]\,.$
This mode and its 68, 95 and 99% central probability intervals are shown in Figure 4, together with the SM and best-fit points. The “scale factor” scenario yields qualitatively the same results. While our results indicate very strong constraints on potential (pseudo)scalar and tensor effects in $sc\nu\ell$ sector of the Weak Effective Theory, they do allow for surprisingly large CP-violating effects in right-handed currents, at the level of $25\%$ of the SM value for the left-handed current. Such a new source of CP violation in the second quark generation would of course be interesting in the context of the observed CP asymmetry in non-leptonic $D$ decays LHCb:2019hro . However, recent analyses of the high $p_{T}$ lepton tails in Drell-Yan processes Fuentes-Martin:2020lea seem to exclude this type of explanation.
Figure 4: Marginalised 1D- and 2D-posterior distributions for the 9 parameters of interest of our WET model in the nominal scenario. As discussed in the text, the imaginary part of $\mathcal{C}_{V_{L}}$ is set to zero using the global unconstrained phase. The “+” and the solid black lines show the SM point $\mathcal{C}_{V,L}=1.01$ and $\mathcal{C}_{i}=$ for all other operators. The “$\times$” and the dashed black lines show the position of our best-fit point. The blue areas are the central 68%, 95%, and 99% integrated probability contours of the posterior distribution obtained from a kernel density estimation.
@@ 4 Conclusion
Our result deviates from the reference values by the Particle Data Group by more than $2.7\,\sigma$. A large part of the observed discrepancy is traced back to a different treatment of the electroweak corrections. Our findings lead to a $4.3\sigma$ and $5.2\sigma$ deviation from unitarity in the second-row and second-column, respectively of the CKM quark-mixing matrix. In light of the observed tension between the lattice QCD results for the $D\to\bar{K}$ form factors, our results should be revised once this tension is clarified.
We set stringent constraints on potential BSM effects in the $sc\nu\ell$ sector of the WET, which limit hypothetical (pseudo)scalar or tensor effects to be below $7\%$ of $\mathcal{C}^{\mathrm{SM}}_{V,L}\simeq 1.01$, the SM value for Wilson coefficient of the left-handed operator. Nevertheless, we find that right-handed currents can still be sizeable. In particular, we find that CP-violating effects in right-handed currents at the level of $\simeq 25\%$ of the SM contribution are not yet excluded. Complementary experimental data, such as the angular distribution of $\Lambda_{c}\to\Lambda(\to p\pi)\ell^{+}\nu$ decays would provide the statistical power to exclude such large CP-violating right-handed contributions.
@@ A.1 Definitions
Note that in this case the two non-vanishing matrix elements are not related by equations of motion and therefore do not share a common decay constant.
The simplest set of three form factors arises in $D\to K\ell^{+}\nu$ decays, from vector, scalar, and tensor currents. A common definition of the form factors reads
@@ A.2 Dispersive Bounds
$\chi^{(t)}_{V}$ $1$ $1.38$ $f_{0}^{D\to\bar{K}}$, $f_{t,V}^{\Lambda_{c}\to\Lambda}$
$\chi^{(0)}_{V}$ $2$ $1.52/m_{c}^{2}$ $f_{+}^{D\to\bar{K}}$, $f_{0,V}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}^{*}\,(2.112,0.274(6))$
$\chi^{(\perp)}_{V}$ $2$ $1.52/m_{c}^{2}$ $f_{\perp,V}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}^{*}\,(2.112,0.275(6))$
$\chi^{(t)}_{A}$ $1$ $2.51$ $f_{t,A}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}\,(1.968,0.2499(5))$
$\chi^{(0)}_{A}$ $2$ $0.98/m_{c}^{2}$ $f_{0,A}^{\Lambda_{c}\to\Lambda}$
$\chi^{(\parallel)}_{A}$ $2$ $0.98/m_{c}^{2}$ $f_{\perp,A}^{\Lambda_{c}\to\Lambda}$
$\chi^{(0)}_{T}$ $3$ $1.12/m_{c}^{2}$ $f_{T}^{D\to K}$, $f_{0,T}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}^{*}\,(2.112,0.28(2))$
$\chi^{(\perp)}_{T}$ $3$ $1.12/m_{c}^{2}$ $f_{\perp,T}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}^{*}\,(2.112,0.28(2))$
$\chi^{(0)}_{AT}$ $3$ $0.88/m_{c}^{2}$ $f_{0,T5}^{\Lambda_{c}\to\Lambda}$
$\chi^{(\parallel)}_{AT}$ $3$ $0.88/m_{c}^{2}$ $f_{\perp,T5}^{\Lambda_{c}\to\Lambda}$
Table 4: List of the minimally-subtracted correlators $\chi^{(\lambda)}_{\Gamma}$ relevant to the hadronic matrix elements used in this analysis. We calculate the numerical values based on formulas up to next-to-leading order in $\alpha_{s}$ and power corrections up order $1/m_{c}^{5}$ (from Ref. Bharucha:2010im ) and next-to-next-to-leading order (NNLO) in $\alpha_{s}$ (from Ref. Grigo:2012ji ). The reference values in the $\overline{\text{MS}}$ scheme for the masses of the $c$-quark and $s$-quark and the strong coupling are $m_{c}(m_{c})=1.275\,\textrm{GeV}$, $m_{s}(m_{c})=112\,\textrm{GeV}$, and $\alpha_{s}(m_{c})=0.3996$. For scale dependent quantities we use $\mu=m_{c}(m_{c})$.
The various hadronic decay constants and form factors discussed in section A.1 are genuine nonperturbative quantities. Nevertheless, perturbation theory can assist in elucidating at least some information about them. This is achieved with the framework of dispersive bounds; see Ref. Caprini:2019osi for a text-book introduction. Here, we use a modified formulation of the dispersive bounds Gubernari:2023puw , compared to common formulations Okubo:1973tj ; Boyd:1997kz ; Caprini:1997mu . Following Ref. Gubernari:2023puw , we begin with defining a suitable two-point correlation function $\Pi^{J}_{\Gamma}(Q^{2})$,
@@
with a decomposition into scalar-valued functions $\Pi^{(\lambda)}$
$\Pi^{\mu\nu}_{\Gamma}(q)$ $\equiv\sum_{\lambda}\epsilon^{\mu}(\lambda)\epsilon^{\nu*}(\lambda)\,\Pi^{(\lambda)}_{\Gamma}(q^{2})\,.$ (39)
Here, $\epsilon^{\mu}(\lambda)$ represent any element of the following basis of Lorentz vector,
$\epsilon_{\mu}(t)$ $=\frac{1}{\sqrt{q^{2}}}q_{\mu}=\frac{1}{\sqrt{q^{2}}}(q_{0},0,0,|\vec{q}\,|)\,,$ $\epsilon_{\mu}(0)$ $=\frac{1}{\sqrt{q^{2}}}(|\vec{q}\,|,0,0,q_{0})\,,$ (40)
$\epsilon_{\mu}(\perp)$ $=(0,-1,0,0)\,,$ $\epsilon_{\mu}(\parallel)$ $=(0,0,-i,0)\,.$
corresponding to one timelike ($\lambda=t$) and three spacelike ($\lambda=0,\parallel,\perp$) polarisation vectors. For our analysis, the relevant currents $J_{\Gamma}$ are
@@
is essential to render OPE results for the correlation function finite. For the vector and axial currents, we use analytic results obtained from a calculation to NNLO in $\alpha_{s}$ Grigo:2012ji together with analytic results for the contributions by $q\bar{q}$, $GG$, and $\bar{q}Gq$ vacuum condensates Bharucha:2010im . For the tensor currents, no NNLO calculation is available. We use the NLO and condensate results obtained in Ref. Bharucha:2010im . The values of $\chi_{\Gamma}^{(\lambda)}$ and the minimal number of subtractions $n$ are compiled in table 4. We find that, although the individual condensate contributions are sizable, they largely cancel when taking their sum.
A hadronic representation of the same correlation functions arises from the computation of the imaginary part of $\Pi^{(\lambda)}$ in terms of hadronic matrix elements
@@
For one-particle bound states $H=\bar{D}_{s}^{(*)}$, the contributions read
@@
where $\lambda(s)\equiv\lambda(M_{D}^{2},M_{K}^{2},s)$ denotes the Källen function. Similar relations exist for all $D\to\bar{K}$ and $\Lambda_{c}\to\Lambda$ form factors. The assignment of individual form factors to the quantities $\chi^{(t)}_{\Gamma}$ is provided in table 4. These relations inspire dispersively-bounded parametrisations Okubo:1973tj ; Boyd:1997kz ; Caprini:1997mu of the hadronic form factors. We apply this framework in the form discussed in Refs. Blake:2022vfl ; Amhis:2022vcd ; Gubernari:2023puw , which improves upon previous works by accounting for the integration domain for the dispersive bound and by splitting the bounds by the helicity as discussed above. The final parametrisation for a form factor $f$ takes the form
@@
where we use the usual conformal map from $q^{2}$ to $z$, outer functions $\phi_{f}$, and Blaschke factors $B(z)$. The function $p_{k}^{(f)}$ are a suitable choice of polynomials of order $k$ that are orthonormal on an $f$-specific arc of the unit circle in $z$ Gubernari:2023puw . The manifest benefit of using this parametrisation is the bounded parameter space $|a_{k}^{(f)}|<1$ for all orders in $k$ and all form factors $f$. Moreover, in a global fit, we impose a strong dispersive bound of the form
@@
where the sum over $f$ iterates over all form factors across processes for a fixed current $\Gamma$ and helicity state $\lambda$, i.e., over all such form factors emerging in either $D\to\bar{K}$ and $\Lambda_{c}\to\Lambda$.

### v2

@@ 3 Methods and Results
nominal SM $60.9$ $51$ $16.1$ $240.3\pm 0.3$
CKM $51.7$ $50$ $40.9$ $251.7\pm 0.3$
WET $48.4$ $42$ $23.1$ $250.3\pm 0.3$
scale factor SM $67.4$ $51$ $6.2$ $232.8\pm 0.3$
CKM $48.2$ $50$ $54.5$ $249.1\pm 0.3$
WET $46.6$ $42$ $29.0$ $248.7\pm 0.3$
@@
Figure 2: Comparison of the predictions in our “nominal” scenario for the various branching ratios for $D_{s}^{+}\to\ell^{+}\nu$, $D^{+(0)}\to K^{0}_{S}(K^{-})\ell^{+}\nu$, and $\Lambda_{c}\to\Lambda\ell^{+}\nu$. The decay $D_{s}^{*+}\to\ell^{+}\nu$ is omitted since the large theory uncertainties make a meaningful visual comparison off the predictions with the measurement impossible. Measurements are shown as black diamonds. Prior predictions labelled as “theory only” are shown as purple coloured circles. Posterior predictions for the two fit models are shown as magenta triangles (SM) and cyan squares (CKM).
@@ 3.1 SM Prior Predictions and Fit
To test a possible issue with the normalisation of the lattice QCD results, we float all hadronic nuisance parameters within the SM fit model (“nominal” scenario) and fit to the full experimental likelihood. We find that we can reduce the experimental $\chi^{2}$ to $60.9$ with a $p$ value of $16.1\%$. This happens at the expense of moving away from the a-priori parameter values by $4.0\,\sigma$. We produce posterior-predictive distributions for the relevant absolute branching fractions, which are shown in fig. 2 as magenta triangles. We can still observe that the posterior predictions systematically overshoot the measurements. However, the tensions with respect to the experimental measurements are visibly reduced. At the level of the variety, number, and accuracy of the hadronic matrix elements used in this analysis, we find compatibility between the hadronic matrix elements and the dispersive bounds.
@@ 3.2 CKM Fit
which agrees with the weighted average of the individual fit results. The central value for the “scale factor” result is higher than the above but compatible at the $1\,\sigma$ level, as shown on the right-hand side of fig. 3. As illustrated in fig. 2, floating the CKM parameter $|V_{cs}|$ results in overall better agreement with the data. Contrary to the SM prior predictions and the SM fit posterior predictions, no systematic shift to larger or smaller values for the branching ratios is visible. This is also reflected in table 1. For the “nominal” scenario, we show that our CKM fit yields a reduction of the $\chi^{2}$ in the best-fit point by $9$ at the expense of one degree of freedom. Using Wilks’ theorem, we therefore obtain a preference for the CKM fit over the SM fit at the $2.7\,\sigma$ level. A Bayesian model comparison yields a Bayes factor of $\simeq 8.9\cdot 10^{4}$, decisively in favour of the CKM fit model. For the “scale factor” scenario, we find a similar preference for the CKM fit model.
@@ 3.3 CKM unitarity
The violation of CKM unitarity of around $3\,\sigma$ by the first-row elements of the CKM matrix has received quite some attention Seng:2018yzq ; Belfatto:2019swo ; Grossman:2019bzp ; Crivellin:2020lzu ; Kirk:2020wdk ; Crivellin:2021njn ; Seng:2021nar ; Cirigliano:2022yyo ; Crivellin:2022rhw ; Cirigliano:2023nol . Here, we perform an alternative test of CKM unitarity by probing the normalisation of both the second row and column;
@@
Our results are given in table 3. Under the assumption of $100\%$ positively correlated uncertainties, we find a deficit toward the expectation of CKM second-row and second-column unitarity at the $4.3\,\sigma$ level and $5.2\,\sigma$ level using the nominal results.
@@
Table 3: Results of the CKM unitarity tests. The last two lines contain the squared sum of the CKM elements of the second row or column, as well as the pull to unity assuming Gaussian uncertainties. The uncertainties on the CKM entries are assumed to be 100% positively correlated, which corresponds to the most conservative scenario (proper estimations would give larger pulls).
This large tension with unitarity again strengthens the case to investigate the normalisation issue that we already pointed out in section 3.1.
@@ 3.4 BSM Interpretation
Our findings so far motivate us to investigate further the allowed parameter space for BSM contribution to $c\to s\ell^{+}\nu$ processes. Lifting the assumption of SM dynamics, we fit the 9 parameters discussed in section 2.4. We find the resulting posterior PDF to be multi-modal and each mode to be distinctly non-Gaussian. The individual modes are related through symmetries of our likelihood and therefore feature the same maximum a-posteriori. Here and in our supplementary material, we provide information on one chosen mode of the posterior, which is defined by $\arg\mathcal{C}^{\ell}_{V,L}=0$ and $\mathcal{C}^{\ell}_{V,L}\simeq 1$. In our nominal scenario, we obtain the following 68% probability intervals
$\operatorname{Re}\mathcal{C}^{\ell}_{V,L}$ $=[\phantom{-}0.957,\phantom{-}1.002]\,,$ (27)
$\operatorname{Re}\mathcal{C}^{\ell}_{V,R}$ $=[-0.026,-0.012]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{V,R}$ $=[-0.225,0.225]\,,$
$\operatorname{Re}\mathcal{C}^{\ell}_{S,L}$ $=[-0.019,\phantom{-}0.014]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{S,L}$ $=[-0.030,0.030]\,,$
$\operatorname{Re}\mathcal{C}^{\ell}_{S,R}$ $=[-0.026,\phantom{-}0.006]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{S,R}$ $=[-0.028,0.028]\,,$
$\operatorname{Re}\mathcal{C}^{\ell}_{T}$ $=[-0.021,\phantom{-}0.046]\,,$ $\operatorname{Im}\mathcal{C}^{\ell}_{T}$ $=[-0.068,0.068]\,.$
This mode and its 68, 95 and 99% central probability intervals are shown in Figure 4, together with the SM and best-fit points. The “scale factor” scenario yields qualitatively the same results. While our results indicate very strong constraints on potential (pseudo)scalar and tensor effects in $sc\nu\ell$ sector of the Weak Effective Theory, they do allow for surprisingly large CP-violating effects in right-handed currents, at the level of $23\%$ of the SM value for the left-handed current. Such a new source of CP violation in the second quark generation would, of course, be interesting in the context of the observed CP asymmetry in non-leptonic $D$ decays LHCb:2019hro . However, recent analyses of the high $p_{T}$ lepton tails in Drell-Yan processes Fuentes-Martin:2020lea seem to exclude this type of explanation.
Figure 4: Marginalised 1D- and 2D-posterior distributions for the 9 parameters of interest of our WET model in the nominal scenario. As discussed in the text, the imaginary part of $\mathcal{C}_{V_{L}}$ is set to zero using the global unconstrained phase. The “+” and the solid black lines show the SM point $\mathcal{C}_{V,L}=1.01$ and $\mathcal{C}_{i}=0$ for all other operators. The “$\times$” and the dashed black lines show the position of our best-fit point. The blue areas are the central 68%, 95%, and 99% integrated probability contours of the posterior distribution obtained from a kernel density estimation.
@@ 4 Conclusion
Our result deviates from the reference values by the Particle Data Group by more than $2.7\,\sigma$. A large part of the observed discrepancy is traced back to a different treatment of the electroweak corrections. Our findings lead to a $4.3\sigma$ and $5.2\sigma$ deviation from unitarity in the second row and second column of the CKM quark-mixing matrix. In light of the observed tension between the lattice QCD results for the $D\to\bar{K}$ form factors, our results should be revised once this tension is clarified.
We set stringent constraints on potential BSM effects in the $sc\nu\ell$ sector of the WET, which limit hypothetical (pseudo)scalar or tensor effects to be below $7\%$ of $\mathcal{C}^{\mathrm{SM}}_{V,L}\simeq 1.01$, the SM value for the Wilson coefficient of the left-handed operator. Nevertheless, we find that right-handed currents can still be sizeable. In particular, we find that CP-violating effects in right-handed currents at the level of $\simeq 23\%$ of the SM contribution are not yet excluded. Complementary experimental data, such as the angular distribution of $\Lambda_{c}\to\Lambda(\to p\pi)\ell^{+}\nu$ decays would provide the statistical power to exclude such large CP-violating right-handed contributions.
@@ A.1 Definitions
Note that in this case, the two non-vanishing matrix elements are not related by equations of motion and, therefore, do not share a common decay constant.
The simplest set of three form factors arises in $D\to K\ell^{+}\nu$ decays from vector, scalar, and tensor currents. A common definition of the form factors reads
@@ A.2 Dispersive Bounds
$\chi^{(J=0)}_{V}$ $1$ $1.38$ $f_{0}^{D\to\bar{K}}$, $f_{t,V}^{\Lambda_{c}\to\Lambda}$
$\chi^{(J=1)}_{V}$ $2$ $1.52/m_{c}^{2}$ $f_{+}^{D\to\bar{K}}$, $f_{0,V}^{\Lambda_{c}\to\Lambda}$, $f_{\perp,V}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}^{*}\,(2.112,0.274(6))$
$\chi^{(J=0)}_{A}$ $1$ $2.51$ $f_{t,A}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}\,(1.968,0.2499(5))$
$\chi^{(J=1)}_{A}$ $2$ $0.98/m_{c}^{2}$ $f_{0,A}^{\Lambda_{c}\to\Lambda}$, $f_{\perp,A}^{\Lambda_{c}\to\Lambda}$
$\chi^{(J=1)}_{T}$ $3$ $1.12/m_{c}^{2}$ $f_{T}^{D\to K}$, $f_{0,T}^{\Lambda_{c}\to\Lambda}$, $f_{\perp,T}^{\Lambda_{c}\to\Lambda}$ $\bar{D}_{s}^{*}\,(2.112,0.28(2))$
$\chi^{(J=1)}_{AT}$ $3$ $0.88/m_{c}^{2}$ $f_{0,T5}^{\Lambda_{c}\to\Lambda}$, $f_{\perp,T5}^{\Lambda_{c}\to\Lambda}$
Table 4: List of the minimally-subtracted correlators $\chi^{(\lambda)}_{\Gamma}$ relevant to the hadronic matrix elements used in this analysis. We calculate the numerical values based on formulas up to next-to-leading order in $\alpha_{s}$ and power corrections up order $1/m_{c}^{5}$ (from Ref. Bharucha:2010im ) and next-to-next-to-leading order (NNLO) in $\alpha_{s}$ (from Ref. Grigo:2012ji ). The reference values in the $\overline{\text{MS}}$ scheme for the masses of the $c$-quark and $s$-quark and the strong coupling are $m_{c}(m_{c})=1.275\,\textrm{GeV}$, $m_{s}(m_{c})=112\,\textrm{GeV}$, and $\alpha_{s}(m_{c})=0.3996$. For scale-dependent quantities, we use $\mu=m_{c}(m_{c})$.
The various hadronic decay constants and form factors discussed in section A.1 are genuine nonperturbative quantities. Nevertheless, perturbation theory can assist in elucidating at least some information about them. This is achieved with the framework of dispersive bounds; see Ref. Caprini:2019osi for a textbook introduction. Here, we use a modified formulation of the dispersive bounds Gubernari:2023puw , compared to common formulations Okubo:1973tj ; Boyd:1997kz ; Caprini:1997mu . Following Ref. Gubernari:2023puw , we begin with defining a suitable two-point correlation function $\Pi^{J}_{\Gamma}(Q^{2})$,
@@
with a decomposition into scalar-valued functions $\Pi^{(i)}$
$\Pi^{\mu\nu}_{\Gamma}(q)$ $\equiv\sum_{i}\mathcal{S}^{\mu\nu}_{i}\,\Pi^{(i)}_{\Gamma}(q^{2})\,.$ (39)
Here, $\mathcal{S}^{\mu\nu}_{i}$ represents one of two structures with definite angular momentum $J$:
$\mathcal{S}^{\mu\nu}_{(J=1)}=\left(\frac{q^{\mu}q^{\nu}}{q^{2}}-g^{\mu\nu}\right)\,,\qquad\mathcal{S}^{\mu\nu}_{(J=0)}=\frac{q^{\mu}q^{\nu}}{q^{2}}$ (40)
For our analysis, the relevant currents $J_{\Gamma}$ are
@@
is essential to render OPE results for the correlation function finite. For the vector and axial currents, we use analytic results obtained from a calculation to NNLO in $\alpha_{s}$ Grigo:2012ji together with analytic results for the contributions by $q\bar{q}$, $GG$, and $\bar{q}Gq$ vacuum condensates Bharucha:2010im . For the tensor currents, no NNLO calculation is available. We use the NLO and condensate results obtained in Ref. Bharucha:2010im . The values of $\chi_{\Gamma}^{(i)}$ and the minimal number of subtractions $n$ are compiled in table 4. We find that, although the individual condensate contributions are sizable, they largely cancel when taking their sum.
A hadronic representation of the same correlation functions arises from the computation of the imaginary part of $\Pi^{(i)}$ in terms of hadronic matrix elements
@@
where $\mathcal{P}_{(J=0)}=\mathcal{S}_{(J=0)}$ and $\mathcal{P}_{(J=1)}=\mathcal{S}_{(J=1)}/3$. For one-particle bound states $H=\bar{D}_{s}^{(*)}$, the contributions read
@@
where $\lambda(s)\equiv\lambda(M_{D}^{2},M_{K}^{2},s)$ denotes the Källen function. Similar relations exist for all $D\to\bar{K}$ and $\Lambda_{c}\to\Lambda$ form factors. The assignment of individual form factors to the quantities $\chi^{(i)}_{\Gamma}$ is provided in table 4. These relations inspire dispersively-bounded parametrisations Okubo:1973tj ; Boyd:1997kz ; Caprini:1997mu of the hadronic form factors. We apply this framework in the form discussed in Refs. Blake:2022vfl ; Amhis:2022vcd ; Gubernari:2023puw , which improves upon previous works by accounting for the integration domain for the dispersive bound and by splitting the bounds by the helicity as discussed above. The final parametrisation for a form factor $f$ takes the form
@@
where we use the usual conformal map from $q^{2}$ to $z$, outer functions $\phi_{f}$, and Blaschke factors $B(z)$. The functions $p_{k}^{(f)}$ are a suitable choice of polynomials of order $k$ that are orthonormal on an $f$-specific arc of the unit circle in $z$ Gubernari:2023puw . The manifest benefit of using this parametrisation is the bounded parameter space $|a_{k}^{(f)}|<1$ for all orders in $k$ and all form factors $f$. Moreover, in a global fit, we impose a strong dispersive bound of the form
@@
where the sum over $f$ iterates over all form factors across processes for a fixed current $\Gamma$ and angular momentum $J$, i.e., over all such form factors emerging in either $D\to\bar{K}$ and $\Lambda_{c}\to\Lambda$.